The length of a rectangle is 3 cm more than 2 times its width. If the area of the rectangle is 93 square centimeters, find the width of the rectangle to the nearest thousandth.
Thank you for the service you offer

Answer by mananth(16949)

let width = x
length = 2x+3
..
ARea = L*W
x(2x+3)=93
2x^2+3x=93
2x^2+3x-93=0
quadratic formula
x1=(-b+sqrt(b^2-4ac))/2a
a=2, b=3, c=-93
x1=(-3+sqrt(9+744))/4
x1= 6.110 cm the width
..
x2=(-3-sqrt(9+744))/4
x2= -7.610 Ignore the negative value